One item in the proposed new union contract allows workers in a shipping department 17 minutes for rest, 6 minutes for personal time, and 11 minutes for delays for each four hours worked.

A time study analyst observed a job that is performed continuously and found an average time of 5.3 minutes per cycle for a worker she rated at 92 percent of the observed time. What standard time is applicable for that operation?

(Keep two decimal places in your answer)
Answer #1

Answer: 5.57 mins

Allowance = 17 +6 + 11 = 34 minutes for each four hours (240 minutes)

Allowance % = 100*(34/240) = 14.167%

Observed time = 5.3 minutes

Rating = 92%

Normal time = Observed time * Rating = 5.3*0.92 = 4.876 mins

Allowance @14.167%

Allowance = Normal time*Allowance % = 4.876*0.14167 = 0.69 mins

Standard time = Normal time + Allowance = 4.876 + 0.69 = 5.566 ~ 5.57 mins